OpenStudy (anonymous):

12x−2=20 log 12x−2=log 20 (x−2)log 12=log 20 (x−2)=log 20/log 12 x=2+log 20/log 12 x=2+log12 20 Someone finish this, need answer rounded to the nearest ten-thousandth, Thanks

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 12^{x−2}=20 }\) like this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x= ?

OpenStudy (solomonzelman):

I gave you the calculator in the previous question.

OpenStudy (anonymous):

dont know how to use it..

OpenStudy (anonymous):

\[x-2 = \frac{ \log(20) }{ \log(12) }\]

OpenStudy (anonymous):

can you just find what x=? appreciate it..

OpenStudy (anonymous):

\[x = 2+\frac{ \log(20) }{ \log(12) }\]

OpenStudy (solomonzelman):

http://www.1728.org/logrithm.htm \(\LARGE\color{blue}{ \bf Log_{~base~}(~log~of) }\) example to use the calc. \(\LARGE\color{blue}{ \bf Log_{2}(5) }\) see where it says "logarithm of" ? is what you put in 5 see where it says "Number base" ? click other, and put in 2

OpenStudy (solomonzelman):

When the base is NOT SPECIFIED, it is 10.

OpenStudy (anonymous):

My nemesis logarithm calculator.

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf x=2+\frac{log20}{log12} }\) \(\LARGE\color{blue}{ \bf x=2+log_{12}20 }\) applying the rule to obtain this, \(\LARGE\color{red}{ \bf \frac{logA}{logB} =Log_{B}A}\)

OpenStudy (anonymous):

i guess you dont understand

OpenStudy (solomonzelman):

who , me ?!?!